OpenStudy (anonymous):

HELP! MEDAL REWARDED!! What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10 (−4, −2) and (2, 5) (−2, −4) and (2, 5) (−2, −4) and (5, 2) No Real Solutions

OpenStudy (anonymous):

with the y=x2+12x+30 did you mean y=x^2+12x+30?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well this is what I got, so..

OpenStudy (anonymous):

What is the solution to the following system of equations? y = x^2 + 10x + 11 y = x^2 + x − 7 (−2, −5) (2, −5) (−2, 5) (2, 5)